OpenStudy (anonymous):

Ninety percent of what number is 45?

OpenStudy (anonymous):

let 90 % of x=45 \[x=45\times \frac{ 100 }{ 90 }=?\]

OpenStudy (anonymous):

To translate the question into an algebraic expression, we write 90% as 0.9 and create the variable x to represent the number in question. 0.9 * x = 45 x = 45 / 0.9 = 50. To verify, 90% of 50 is 45.